OpenStudy (ilovecomputers1):

What is the axis of symmetry for f(x) = 2x2 − 4x + 5?

OpenStudy (freckles):

The axis of symmetry is x=-b/(2a)

OpenStudy (freckles):

Determine what a and b are from y=ax^2+bx+c

OpenStudy (ilovecomputers1):

A is 2, B is -4.

OpenStudy (ilovecomputers1):

So it would be x=4/4=1. So that's it right?

OpenStudy (kainui):

@freckles where did you get this random formula from? You should explain things, because this formula did not just come out of thin air. If you yourself do not understand it, then you should strive to learn why, because @ILoveComputers1 is not learning anything here.

OpenStudy (ilovecomputers1):

Yeah I was wondering that as well @Kainui. Thank you.

OpenStudy (freckles):

Yep x=1. You can't just say one. Make sure you write x=1. This represents a vertical line. It is called the axis of symmetry because you can fold your parabola over that line and looking through the paper you will just that curve laying top of the other part of the curve perfectly.

OpenStudy (ilovecomputers1):

I think I get it now. Thank you so much @freckles.

OpenStudy (freckles):

You can add what you like Kaini. Just doing the best I can on an Android.

OpenStudy (kainui):

 I'm bad at memorizing formulas, but we can always figure it out like this. Here I labeled "a" as the x value for symmetry and then we can see that f(x+a) is always the same as f(a-x) so we can set those equal and solve for a that way. But @freckles is not wrong and I think he did a good job of explaining it for that method as well. =)